Circular No. 8428

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 V2 (HARTLEY-IRAS) R. H. McNaught, Siding Spring Observatory, reports his recovery of

comet P/1983 V1 (= 1983v = 1984 III) on CCD images taken with the 1.0-m f/8 reflector on Nov. 3.7 and 5.8 UT. The comet appears slightly diffuse with no tail. The indicated correction to the predictions on MPC 45657 (ephemeris on MPC 52656) and in the 2004 Comet Handbook is ΔT ≃ −4.8 days. The available 2004 astrometry, along with the following orbital elements and an ephemeris, appear on MPEC 2004-V25.

Epoch = 2005 July 9.0 TT

T = 2005 June 20.8367 TT  ω = 47.^{o}0753
e = 0.835083  Ω = 1.4002 }2000.0
q = 1.275129 AU  i = 95.6969

a = 7.731949 AU n^{o} = 0.0458427 P = 21.500 years

SUPERNOVA 2004fu IN NGC 6949 Further to IAUC 8406, R. Arbour reports his discovery of an apparent

supernova on unfiltered CCD images taken with a 0.3-m f/6.3 Schmidt-Cassegrain reflector on Nov. 4.874 (at mag ≈ 16.8) and 5.813 (mag ≈ 16.5). The new object is located at α = 20^{h}35^{m}11^{s}.54, δ = +64^{o}48′25′′.7 (equinox 2000.0), which is 29′′ east and 15′′ north of the center of NGC 6949. An image taken by D. Briggs with a 0.60-m f/4.7 reflector on Nov. 5.864 yields position end figures 11^{s}.62, 26′′.5. Nothing is visible at this location on a patrol image taken by Arbour on Oct. 31.907 (limiting mag ≈ 17.5) or on Digitized Sky Survey images from 1991 (limiting blue mag 19.0) and 1992 (limiting red mag 19.0).

SUPERNOVA 2004ff IN ESO 552-G40 M. Modjaz, P. Challis, R. Kirshner, A. Garg, and C. Stubbs, Harvard-

Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 390–900 nm) of SN 2004ff (cf. IAUC 8425), obtained by Challis, Garg, and Stubbs on Nov. 3.30 UT with the Magellan Clay 6.5-m Telescope (+ LDSS-2), shows it to be a supernova of type Ic. The spectrum is similar to spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11) taken a few days after maximum.

2004 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green